# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

### v1

@@ 4.1 Multi-Modal Adapter Architecture
We build on the adapter-based design introduced in [38] to incorporate a lightweight and efficient tuning mechanism for vision-language models. This architecture has proven effective in few-shot generalization settings, where pre-trained CLIP models are fine-tuned on a limited number of base classes and tested on base and novel, unseen categories.
The motivation for this design stems from two empirical findings: (i) higher layers of both image and text encoders in CLIP contain more discriminative and dataset-specific features, while lower layers preserve general, transferable knowledge; and (ii) larger modality gaps between text and image encoders are observed in the lower layers, making cross-modal alignment particularly challenging in the early stages of the network. [38].
@@
Each adapter consists of:
  • A down-projection layer that reduces the input dimension,
  • A shared projection layer that facilitates interaction between the modalities,
  • An up-projection layer that restores the original dimension.
This three-part structure allows the adapter to first transform features into a low-dimensional space, fuse them through a shared module, and then project them back. Formally, for the visual adapter at the $j$-th block:
$\mathcal{A}_{j}^{(I)}(\boldsymbol{z}_{j}^{(I)})=W_{ju}^{(I)}\cdot\delta(W_{js}\cdot\delta(W_{jd}^{(I)}\cdot\boldsymbol{z}_{j}^{(I)})),\quad j\in\{\ell,\cdots,L\},$ (7)
and similarly for the textual adapter:
$\mathcal{A}_{j}^{(T)}(\boldsymbol{z}_{j}^{(T)})=W_{ju}^{(T)}\cdot\delta(W_{js}\cdot\delta(W_{jd}^{(T)}\cdot\boldsymbol{z}_{j}^{(T)})),\quad j\in\{\ell,\cdots,L\},$ (8)
where $\boldsymbol{z}_{j}^{(I)}$ and $\boldsymbol{z}_{j}^{(T)}$ denote the input hidden states at the $j$-th transformer layer for the vision and text encoders, $W_{jd}^{(I)}$ and $W_{jd}^{(T)}$ are the down-projection matrices, $W_{js}$ is the shared projection matrix used across both modalities, $W_{ju}^{(I)}$ and $W_{ju}^{(T)}$ are the up-projection matrices, and $\delta(\cdot)$ denotes the non-linear activation function (e.g., GELU), applied element-wise.
@@
In contrast to methods that inject prompts or adapters across all layers [12, 78, 34] or some lower layers [80, 73, 61, 62], this selective, top-layer insertion strategy reduces the number of trainable parameters while maintaining cross-modal adaptability.
@@ 4.2 Generalization and Personalization via pFedMMA
This selective update mechanism provides several key benefits:
  • Local personalization: By allowing clients to optimize their own up- and down-projection matrices, each client can adapt the representation space to their unique local data distribution. This is particularly effective under label and feature heterogeneity.
  • Global generalization: The shared projection matrix is collaboratively trained across clients and is responsible for aligning the modalities in a consistent global space. This facilitates transferability and enables the model to generalize well across diverse domains and tasks.
  • Communication efficiency: Since the shared projection layer is low-dimensional compared to the full model or full adapter stack, transmitting only the shared component during communication rounds results in significantly reduced communication cost.
@@
$W\in\left\{W_{jd,i}^{(I)},W_{ju,i}^{(I)},W_{jd,i}^{(T)},W_{ju,i}^{(T)},W_{js,i}\right\},\quad j\in\{\ell,\cdots,L\},\quad i\in\{1,\cdots,N\},$ (9)
are updated for $E$ local epochs using gradient descent:
$W^{t,e}=W^{t,e-1}-\eta\nabla\mathcal{L}_{ce}(W^{t,e-1}),$ (10)
where $\eta$ is the learning rate, $\mathcal{L}_{ce}$ is the cross-entropy loss, and $e\in\{1,\cdots,E\}$.
After local updates, only the shared projection parameters $W_{js,i}^{t,E}$ are uploaded to the server. These are aggregated across all participating clients to obtain the updated global shared adapter:
$W_{js}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}W_{js,i}^{t,E},$ (11)
where $N$ is the number of participating clients. In contrast, the up- and down-projection parameters remain local and are not shared or averaged.
@@ 5.1 Experimental Setup
Datasets and Data Heterogeneity. To evaluate the effectiveness of pFedMMA, we conduct experiments across nine public benchmark datasets that cover various types of data heterogeneity, including label shift, feature shift, and domain shift. Following prior work such as [23], we use six visual classification datasets: SUN397 [81], Flowers102 [82], DTD [83], Caltech101 [84], UCF101 [85], and Food101 [86]. We refer to these as the CLIP datasets. To simulate severe label heterogeneity, we apply a pathological non-IID setting in which each client is assigned a distinct, non-overlapping set of classes. Clients are trained on their local classes and evaluated on their own classes, on base classes held by other clients, and on novel classes that are unseen during the training process.
@@
To evaluate performance under feature and domain shift, we utilize two widely adopted multi-domain datasets: DomainNet [87], which consists of six distinct domains, and Office-Caltech10 [88], which includes four domains. Following prior studies, each client is assigned data from a single domain, ensuring that every domain is represented by a group of clients in the federation. To introduce additional heterogeneity and simulate realistic federated learning scenarios, we further partition the data within each domain using a symmetric Dirichlet distribution with $\beta=0.3$. This setup introduces both feature shift across domains and label shift within domains. All domains participate in both training and evaluation phases, allowing us to assess cross-domain generalization and personalization performance in more realistic federated conditions.
@@
For personalization evaluation, we include CIFAR-10 [89] and CIFAR-100 [90]. These datasets are partitioned among clients using a symmetric Dirichlet distribution with concentration parameter $\beta=0.3$, which creates varying degrees of label skew across clients. Additionally, we apply the same pathological class split as used in the CLIP datasets to test robustness under extreme heterogeneity. Further details on the dataset configurations and partitioning strategies can be found in Appendix.
@@
Baselines. We evaluate pFedMMA across all experimental settings, including generalization, personalization, and domain generalization, using a consistent set of five baselines. Zero-shot CLIP [1] serves as a non-adaptive reference model that uses fixed hand-crafted prompt templates such as "a photo of a [class]" without any task-specific learning. PromptFL [23] represents a standard federated prompt learning approach in which a shared prompt is collaboratively learned across clients using FedAvg. FedPGP [29] introduces prompt-wise contrastive learning to encourage consistency between global and local prompts. FedOTP [28] applies unbalanced Optimal Transport to align global knowledge with client-specific prompt representations. Finally, pFedMoAP [30] leverages a Mixture-of-Experts design that enables each client to access both local and non-local prompt experts through a lightweight attention-based gating mechanism. These baselines cover a diverse range of federated adaptation strategies, providing a strong benchmark for assessing the performance of pFedMMA across different types of heterogeneity.
Implimentation Details. All methods, including pFedMMA and all baselines, are implemented on top of a frozen CLIP model. We use two backbone architectures, ResNet50 [35] and ViT-B16 [36], and default to ViT-B16 unless otherwise specified. For the CIFAR-10 and CIFAR-100 datasets, we simulate a large-scale federated environment with $100$ clients, using a Dirichlet distribution with $\beta=0.3$ and a $10$ percent client participation rate per communication round. Training runs for $50$ rounds with 1 local epoch per round. For the CLIP datasets, each is split into $10$ clients with non-overlapping classes, using $100$ percent participation, $2$ local epochs, and $50$ communication rounds. In the case of DomainNet and Office-Caltech10, each domain of these two datasets is partitioned to $2$ clients with with Dirichlet $\beta=0.3$, resulting in $N=12$ for DomainNet and $N=8$ for Office-Caltech10. We use SGD with a learning rate of $0.001$, and batch sizes of $32$ for training and $100$ for testing. In pFedMMA, adapter modules are implemented with a bottleneck dimension of $32$, and the scaling factor for adapter outputs is set to $0.001$. Further implementation details are provided in Appendix.
Table 1: Detailed results for the 7 datasets with ViT-B/16 as backbone. Top-1 accuracy averaged over 3 random seeds is reported. Highest value is highlighted in bold, and the second highest is underlined.
@@
CLIP [1] $76.36$ $\bm{76.81}$ $\bm{81.21}$ $78.03$ $69.41$ $69.38$ $\bm{75.52}$ $71.32$ $67.89$ $\underline{69.23}$ $\bm{76.88}$ $71.12$ $54.26$ $54.86$ $\bm{59.18}$ $56.02$
PromptFL [23] $89.92$ $84.04$ $75.54$ $\underline{82.15}$ $78.50$ $\underline{74.85}$ $69.21$ $\underline{73.99}$ $97.17$ $\bm{76.19}$ $66.31$ $\bm{77.93}$ $83.69$ $\bm{77.31}$ $48.37$ $\underline{65.85}$
FedPGP [29] $95.38$ $\underline{76.49}$ $71.68$ $79.10$ $94.29$ $54.88$ $57.76$ $65.02$ $\underline{99.67}$ $72.44$ $58.65$ $73.37$ $89.03$ $\underline{71.03}$ $50.94$ $\bm{66.75}$
FedOTP [28] $\underline{97.34}$ $17.99$ $36.69$ $31.08$ $\underline{94.50}$ $11.51$ $14.86$ $18.21$ $99.65$ $14.62$ $30.49$ $26.97$ $\bm{98.08}$ $20.79$ $35.36$ $34.65$
pFedMoAP [30] $\bm{97.90}$ $61.82$ $66.60$ $71.05$ $\bm{95.93}$ $31.18$ $35.40$ $42.41$ $\bm{99.81}$ $43.70$ $48.37$ $55.99$ $\underline{96.43}$ $53.60$ $48.21$ $60.28$
pFedMMA (Ours) $94.00$ $\underline{76.49}$ $\underline{81.07}$ $\bm{82.85}$ $91.72$ $\bm{69.56}$ $\underline{75.41}$ $\bm{77.84}$ $90.22$ $68.66$ $\underline{76.74}$ $\underline{77.56}$ $85.69$ $55.38$ $\underline{59.12}$ $64.32$
@@
CLIP [1] $89.45$ $89.42$ $\underline{96.81}$ $91.77$ $96.14$ $97.22$ $\bm{94.21}$ $95.84$ $89.40$ $\underline{89.42}$ $\bm{90.70}$ $89.84$ $68.00$ $\underline{68.15}$ $\bm{75.18}$ $70.29$
PromptFL [23] $96.82$ $\underline{93.06}$ $96.39$ $\underline{95.39}$ $97.11$ $\bm{97.70}$ $92.46$ $\underline{95.70}$ $91.28$ $\bm{89.93}$ $90.28$ $\underline{90.49}$ $84.88$ $\bm{79.21}$ $65.73$ $\underline{75.72}$
FedPGP [29] $96.62$ $\bm{95.17}$ $\bm{97.15}$ $\bm{96.31}$ $99.42$ $94.94$ $90.88$ $94.95$ $93.70$ $86.38$ $87.14$ $88.96$ $\underline{94.94}$ $60.62$ $59.26$ $68.33$
FedOTP [28] $\bm{100.00}$ $11.60$ $51.22$ $25.92$ $\underline{99.94}$ $36.47$ $62.77$ $56.22$ $95.69$ $17.29$ $37.97$ $31.70$ $93.54$ $13.62$ $24.19$ $23.91$
pFedMoAP [30] $\underline{99.92}$ $77.61$ $92.05$ $88.87$ $99.92$ $94.07$ $92.43$ $95.37$ $\bm{97.49}$ $69.86$ $83.51$ $82.09$ $\bm{95.79}$ $62.74$ $66.23$ $72.33$
pFedMMA (Ours) $99.85$ $87.00$ $96.48$ $93.40$ $\bm{99.95}$ $\underline{97.26}$ $\underline{94.16}$ $\bm{97.07}$ $\underline{96.95}$ $89.40$ $\underline{90.67}$ $\bm{92.22}$ $93.66$ $68.13$ $\underline{74.91}$ $\bm{77.51}$
Table 2: Results on DomainNet and Office-Caltech10 with feature shift and label shift with $\beta=0.3$ partition into $2$ clients/domain.
Datasets DomainNet Office-Caltech10
@@
Domains Clipart Infograph Painting Quickdraw Real Sketch Amazon Caltech DSLR Webcam Avg.
CLIP [1] $11.82$ $9.23$ $9.45$ $10.96$ $10.52$ $9.79$ $11.78$ $6.21$ $9.92$ $8.51$ $9.82$
PromptFL [23] $11.55$ $12.80$ $13.40$ $8.67$ $20.16$ $5.26$ $10.50$ $15.36$ $\bm{25.95}$ $3.45$ $12.63$
FedPGP [29] $25.61$ $\underline{33.52}$ $\bm{26.51}$ $\underline{23.62}$ $23.62$ $\underline{24.89}$ $21.02$ $\underline{21.38}$ $12.70$ $\underline{22.07}$ $23.49$
pFedMoAP [30] $\underline{25.42}$ $30.15$ $\underline{22.91}$ $20.83$ $\underline{24.85}$ $24.59$ $\bm{24.01}$ $19.13$ $\underline{25.40}$ $\bm{23.68}$ $24.10$
pFedMMA (Ours) $\bm{26.61}$ $\bm{44.42}$ $16.94$ $\bm{46.09}$ $\bm{54.11}$ $\bm{37.46}$ $\underline{23.73}$ $\bm{25.88}$ $23.67$ $15.00$ $\bm{31.49}$
Table 3: Accuracy comparison (%) on the Dirichlet Non-IID setting in CIFAR-10 and CIFAR-100 over 100 clients.
Methods CIFAR-10 CIFAR-100
CLIP [1] 87.95 64.90
PromptFL [23] 92.95 73.48
FedPGP [29] 92.11 72.89
FedOTP [28] 95.82 73.83
pFedMoAP [30] 95.92 75.70
pFedMMA (Ours) 96.79 78.33

### v2

@@ 3.1 Multi-Modal Adapter Architecture
We build on the adapter-based design introduced in Yang et al. (2024) to incorporate a lightweight and efficient tuning mechanism for vision-language models. This architecture has proven effective in few-shot generalization settings, where pre-trained CLIP models are fine-tuned on a limited number of base classes and tested on base and novel, unseen categories.
The motivation for this design stems from two empirical findings: (i) higher layers of both image and text encoders in CLIP contain more discriminative and dataset-specific features, while lower layers preserve general, transferable knowledge; and (ii) larger modality gaps between text and image encoders are observed in the lower layers, making cross-modal alignment particularly challenging in the early stages of the network. Yang et al. (2024).
@@
Each adapter consists of: (i) A down-projection layer that reduces the input dimension, (ii) A shared projection layer that facilitates interaction between the modalities, (iii) An up-projection layer that restores the original dimension.
This three-part structure allows the adapter to first transform features into a low-dimensional space, fuse them through a shared module, and then project them back. Formally, for the visual adapter (indexed by $(I)$) and the textual adapter (indexed by $(T)$) at the $j$-th block:
$\mathcal{A}_{j}^{(o)}({\bm{z}}_{j}^{(o)})={\bm{W}}_{ju}^{(o)}\cdot\delta({\bm{W}}_{js}\cdot\delta({\bm{W}}_{jd}^{(o)}\cdot{\bm{z}}_{j}^{(o)})),\quad o\in\{I,T\},\quad j\in\{\ell,\cdots,L\},$ (2)
where ${\bm{z}}_{j}^{(I)}$ and ${\bm{z}}_{j}^{(T)}$ denote the input hidden states at the $j$-th transformer layer for the vision and text encoders, ${\bm{W}}_{jd}^{(I)}$ and ${\bm{W}}_{jd}^{(T)}$ are the down-projection matrices, ${\bm{W}}_{js}$ is the shared projection matrix used across both modalities, ${\bm{W}}_{ju}^{(I)}$ and ${\bm{W}}_{ju}^{(T)}$ are the up-projection matrices, and $\delta(\cdot)$ denotes the non-linear activation function (e.g., GELU), applied element-wise.
@@
In contrast to methods that inject prompts or adapters across all layers Chen et al. (2022); Houlsby et al. (2019); Hu et al. (2021) or some lower layers Khattak et al. (2023a; b); Zhou et al. (2022b; a), this selective, top-layer insertion strategy reduces the number of trainable parameters while maintaining cross-modal adaptability.
@@ 3.2 Generalization and Personalization via pFedMMA
This selective update mechanism provides several key benefits: (i) Local personalization: By allowing clients to optimize their own up- and down-projection matrices, each client can adapt the representation space to their unique local data distribution. This is particularly effective under label and feature heterogeneity. (ii) Global generalization: The shared projection matrix is collaboratively trained across clients and is responsible for aligning the modalities in a consistent global space. This facilitates transferability and enables the model to generalize well across diverse domains and tasks. (iii) Communication efficiency: Since the shared projection layer is low-dimensional compared to the full model or full adapter stack, transmitting only the shared component during communication rounds results in significantly reduced communication cost. Extensive communication and computational cost analysis are provided in Appendix E.
@@
${\bm{W}}\in\left\{{\bm{W}}_{jd,i}^{(I)},{\bm{W}}_{ju,i}^{(I)},{\bm{W}}_{jd,i}^{(T)},{\bm{W}}_{ju,i}^{(T)},{\bm{W}}_{js,i}\right\},\quad j\in\{\ell,\cdots,L\},\quad i\in\{1,\cdots,N\},$ (3)
are updated for $E$ local epochs using gradient descent: ${\bm{W}}^{t,e}_{i}={\bm{W}}^{t,e-1}_{i}-\eta\nabla\mathcal{L}_{ce}({\bm{W}}^{t,e-1}_{i}),$ where $\eta$ is the learning rate, $\mathcal{L}_{ce}$ is the cross-entropy loss, and $e\in\{1,\cdots,E\}$.
After local updates, only the shared projection parameters ${\bm{W}}_{js,i}^{t,E}$ are uploaded to the server. These are aggregated across all participating clients to obtain the updated global shared adapter: ${\bm{W}}_{js}^{t+1}=\frac{1}{N}\sum_{i=1}^{N}{\bm{W}}_{js,i}^{t,E},$ where $N$ is the number of participating clients. In contrast, the up- and down-projection parameters remain local and are not shared or averaged.
@@ 4.1 Experimental Setup
Datasets and Data Heterogeneity. To evaluate the effectiveness of pFedMMA, we conduct experiments across eleven public benchmark datasets that cover various types of data heterogeneity, including label shift and feature shift. Following prior work such as Guo et al. (2023b), we use seven visual classification datasets: SUN397 Xiao et al. (2010), OxfordPets Parkhi et al. (2012), Flowers102 Nilsback & Zisserman (2008), DTD Cimpoi et al. (2014), Caltech101 Fei-Fei et al. (2004), UCF101 Soomro et al. (2012), and Food101 Bossard et al. (2014). We refer to these as the CLIP datasets. To simulate severe label heterogeneity, we apply a pathological non-IID setting in which each client is assigned a distinct, non-overlapping set of classes. Clients are trained on their local classes and evaluated on their own classes, on base classes held by other clients, and on novel classes that are unseen during the training process.
@@
To evaluate performance under feature shift, we utilize two widely adopted multi-domain datasets: DomainNet Peng et al. (2019), which consists of six distinct domains, and Office-Caltech10 Gong et al. (2012), which includes four domains. Following prior studies, each client is assigned data from a single domain, ensuring that every domain is represented by a group of clients in the federation. To introduce additional heterogeneity and simulate realistic federated learning scenarios, we further partition the data within each domain using a symmetric Dirichlet distribution with concentration parameter $\beta$. This setup introduces both feature shift across domains and label shift within domains. All domains participate in both training and evaluation phases, allowing us to assess cross-domain generalization and personalization performance in more realistic federated conditions.
@@
For personalization evaluation, we include CIFAR-10 Krizhevsky et al. (2010) and CIFAR-100 Krizhevsky et al. (2009). These datasets are partitioned among clients using a Dirichlet distribution, which creates varying degrees of label skew across clients. Additionally, we apply the same pathological class split as used in the CLIP datasets to test robustness under extreme heterogeneity. Further details on the dataset configurations and partitioning strategies can be found in Appendix C.1.
@@
Baselines. We evaluate pFedMMA across all experimental settings, including generalization, personalization, and domain generalization, using a consistent set of five baselines. Zero-shot CLIP Radford et al. (2021) serves as a non-adaptive reference model that uses fixed hand-crafted prompt templates such as ”a photo of a [class]” without any task-specific learning. PromptFL Guo et al. (2023b) represents a standard federated prompt learning approach in which a shared prompt is collaboratively learned across clients using FedAvg. FedPGP Cui et al. introduces prompt-wise contrastive learning to encourage consistency between global and local prompts. FedOTP Li et al. (2024) applies unbalanced Optimal Transport to align global knowledge with client-specific prompt representations. Finally, pFedMoAP Luo et al. (2025) leverages a Mixture-of-Experts design that enables each client to access both local and non-local prompt experts through a lightweight attention-based gating mechanism. These baselines cover a diverse range of federated adaptation strategies, providing a strong benchmark for assessing the performance of pFedMMA across different types of heterogeneity.
Table 1: Top-1 accuracy (%) of different methods across 7 datasets in the 16-shot setting.
@@
CLIP Radford et al. (2021) 76.36 76.81 81.21 78.03 $69.41$ $69.38$ $75.52$ $71.32$ 67.89 69.23 76.88 71.12 $54.26$ $54.86$ $59.18$ $56.02$
PromptFL Guo et al. (2023b) 88.93 88.95 75.36 83.09 $77.73$ $77.71$ $72.96$ $76.07$ 97.37 97.06 63.62 82.66 $80.23$ $80.21$ $45.29$ $63.81$
FedPGP Cui et al. 95.38 76.49 71.68 79.09 $94.29$ $54.88$ $57.76$ $65.02$ 99.67 72.44 58.65 73.37 $89.03$ $71.03$ $50.94$ $66.75$
FedOTP Li et al. (2024) 97.34 18.00 36.69 31.08 $94.50$ $11.51$ $14.86$ $18.21$ 99.65 14.62 30.49 26.97 $98.08$ $20.79$ $35.36$ $34.65$
pFedMoAP Luo et al. (2025) 97.89 61.82 66.60 71.05 $95.93$ $31.18$ $35.40$ $42.41$ 99.81 43.70 48.37 55.99 $96.43$ $53.60$ $48.21$ $60.28$
pFedMMA (Ours) 97.17 77.40 81.49 84.15 $94.06$ $70.99$ $76.37$ 79.34 $95.58$ $71.54$ $76.00$ 79.79 $97.45$ $55.44$ $61.55$ 67.35
$\Delta$ 0.74-0.74%  +1.191.19%  +13.6913.69%  +6.46.4% 1.95-1.95%  +29.3529.35%  +32.2232.22%  +22.0222.02% 4.24-4.24% 1.24-1.24%  +29.5829.58%  +8.758.75% 0.64-0.64% 21.95-21.95%  +20.8320.83%  +0.90.9%
@@
CLIP Radford et al. (2021) $89.45$ $89.42$ $96.81$ $91.77$ $96.14$ $97.22$ $94.21$ $95.84$ $89.40$ $89.42$ $90.70$ $89.84$ $68.00$ $68.15$ $75.18$ $70.29$
PromptFL Guo et al. (2023b) $96.35$ $96.28$ $97.26$ $96.63$ $97.77$ $98.19$ $92.58$ $96.11$ $90.48$ $90.50$ $91.37$ $90.78$ $82.57$ $82.73$ $64.47$ $75.55$
FedPGP Cui et al. $96.62$ $95.17$ $97.15$ $96.31$ $99.42$ $94.94$ $90.88$ $94.95$ $93.70$ $86.38$ $87.14$ $88.96$ $94.94$ $60.62$ $59.26$ $68.33$
FedOTP Li et al. (2024) $100.00$ $11.60$ $51.22$ $25.92$ $99.94$ $36.47$ $62.77$ $56.23$ $95.69$ $17.29$ $37.97$ $31.70$ $93.54$ $13.62$ $24.19$ $23.91$
pFedMoAP Luo et al. (2025) $99.92$ $77.61$ $92.05$ $88.87$ $99.92$ $94.07$ $92.43$ $95.37$ $97.49$ $69.86$ $83.51$ $82.09$ $95.79$ $62.74$ $66.23$ $72.33$
pFedMMA (Ours) $100.00$ $88.50$ $96.60$ 94.78 $100.00$ $96.53$ $94.29$ 96.88 $97.45$ $89.15$ $90.77$ 92.32 $95.63$ $69.61$ $74.88$ 78.58
$\Delta$ 0% 7.01-7.01% 0.57-0.57% 1.59-1.59%  +0.060.06%  +1.671.67%  +2.012.01%  +1.581.58% 0.04-0.04%  +3.213.21%  +4.174.17%  +3.783.78% 0.17-0.17%  +10.9510.95%  +13.0613.06%  +8.648.64%
Figure 3: Local and harmonic mean (HM) accuracies of various methods across different shot settings.
Implimentation Details. All methods, including pFedMMA and all baselines, are implemented on top of a frozen CLIP model. We use two backbone architectures, ViT-B16 and ViT-B32 Dosovitskiy et al. , and default to ViT-B16 unless otherwise specified. For the CLIP datasets, each is split into $10$ clients with non-overlapping classes, using $100$ percent participation, $2$ local epochs, and $50$ communication rounds. For the CIFAR-10 and CIFAR-100 datasets, we simulate a large-scale federated environment with $100$ clients, using a varying Dirichlet distribution and a $10$ percent client participation rate per communication round. Training runs for $50$ rounds with 1 local epoch per round. In the case of DomainNet and Office-Caltech10, each domain of these two datasets is partitioned to $1/2$ clients, resulting in $N=6/12$ for DomainNet and $N=4/8$ for Office-Caltech10. We use SGD with a learning rate of $0.001$, and batch sizes of $32$ for training and $100$ for testing. Further implementation details are provided in Appendix.
Table 2: Accuracy comparison (%) on the Dirichlet Non-IID setting in CIFAR-10 and CIFAR-100.
Dataset CIFAR-100 CIFAR-10
$\beta$ 0.1 0.3 0.5 1 5 10 0.1 0.3 0.5 1 5 10
CLIP Radford et al. (2021) 64.93 64.90 65.00 64.95 64.94 64.91 87.98 87.95 87.93 87.98 88.02 87.98
PromptFL Guo et al. (2023b) 75.34 73.48 72.85 72.83 72.21 72.41 92.80 92.95 94.34 93.89 93.31 93.02
FedPGP Cui et al. 74.72 72.89 74.85 74.18 74.07 73.90 91.69 93.19 93.21 92.98 93.04 92.91
FedOTP Li et al. (2024) 77.53 73.83 72.21 70.99 69.40 68.97 97.23 95.82 94.64 93.10 91.87 91.67
pFedMoAP Luo et al. (2025) 80.29 75.70 75.68 74.53 73.00 72.61 97.13 95.92 94.86 93.97 92.67 92.65
pFedMMA (Ours) 81.82 78.33 76.92 75.70 74.03 73.65 97.37 96.92 95.82 94.82 93.52 93.07
$\Delta$  +1.911.91%  +3.473.47%  +1.641.64%  +1.571.57% 0.05-0.05% 0.34-0.34%  +0.140.14%  +1.041.04%  +1.011.01%  +0.90.9%  +0.230.23%  +0.050.05%
Table 3: Test accuracy (%) of different methods on DomainNet and Office-Caltech10 with lable shift and domain shift using Dirichlet partitioning ($\beta=0.5$).
DomainNet Office-Caltech10
@@
Method Clipart Infograph Painting Quickdraw Real Sketch Avg. Amazon Caltech DSLR Webcam Avg.
CLIP Radford et al. (2021) 8.99 10.69 11.20 10.85 9.53 9.39 10.11 11.78 6.21 9.92 8.51 9.11
PromptFL Guo et al. (2023b) 11.02 1.65 11.20 8.95 13.89 20.75 11.24 10.35 15.83 32.06 7.13 16.34
FedPGP Cui et al. 24.77 31.87 23.87 22.87 22.40 23.64 24.90 20.34 19.12 20.85 22.52 20.71
pFedMoAP Luo et al. (2025) 24.77 30.93 26.09 20.46 22.59 23.10 24.65 20.01 24.45 18.02 15.73 19.55
pFedMMA (Ours) 50.38 23.81 60.27 61.44 40.35 46.79 47.17 9.26 29.15 33.26 13.64 21.33
Table 4: Average performance across six datasets using the ViT-B/32 backbone under different shot settings (1, 2, 4, 8, and 16).
Method 1 Shot 2 Shots 4 Shots 8 Shots 16 Shots
Local Base Novel HM Local Base Novel HM Local Base Novel HM Local Base Novel HM Local Base Novel HM
FedPGP 79.31 79.35 80.25 79.53 82.29 81.43 77.08 80.07 85.86 81.91 74.50 79.68 89.07 80.33 73.61 79.63 93.57 70.79 68.89 74.66
FedOTP 86.84 11.20 20.66 19.50 88.69 11.29 23.12 20.04 91.51 11.02 21.19 19.33 92.44 9.95 19.39 17.25 94.60 9.97 16.74 16.46
pFedMoAP 93.18 45.53 54.08 56.99 95.05 47.14 58.33 59.37 96.53 44.14 54.34 56.26 96.88 42.53 50.93 53.83 97.21 29.83 44.15 44.86
pFedMMA (Ours) 82.14 76.78 79.35 79.31 84.40 77.04 79.34 80.09 86.82 76.85 79.32 80.67 88.07 76.79 79.01 80.83 90.02 76.57 79.75 81.29
$\Delta$ 0.28-0.28%  +0.020.02%  +1.241.24%  +1.511.51%  +8.888.88%
Table 5: Ablation study on pFedMMA design choices, including scaling factor, adapter dimension, starting layer, and adapter sharing strategies.
$\alpha$ Local Base Novel HM
0.0001 91.40 5.06 5.76 7.62
0.0005 91.45 46.34 59.16 59.15
0.001 90.91 72.21 78.48 79.37
0.005 91.03 78.65 81.77 83.25
0.01 91.47 78.12 81.56 82.03
$\ell\to L$ Local Base Novel HM
12 96.49 76.85 81.37 83.57
10 $\to$ 12 96.61 78.14 81.98 84.38
8 $\to$ 12 95.75 78.43 81.82 84.27
6 $\to$ 12 91.53 78.53 81.76 83.32
5 $\to$ 12 91.58 78.67 81.71 83.38
Dims Local Base Novel HM
8 89.15 72.23 76.41 78.24
16 89.84 72.51 77.36 78.93
32 90.91 72.21 78.33 79.37
64 91.55 71.37 78.31 79.15
128 91.78 72.55 78.23 79.68
Method DTD Caltech Flowers OxfordPets
Baseline 1 61.10 96.61 73.82 91.91
Baseline 2 62.19 98.14 77.04 92.75
pFedMMA 76.38 99.48 86.34 97.39